Prove that if a complex number z is an nth root of a real number a , then so
is z bar.
1
Expert's answer
2020-02-20T09:06:40-0500

We know that for every complex number which is in the form of a+ib

now,

for any positive integer n let take nth root in the form of x+iy

so,(a+ib)n = x+iy

Now use De Moivre's Theorem to extend it

so complex number

"z=cos\theta + isin \theta"

which has n distinct root can be determined by

"\sqrt [n]{z}=\sqrt[n]{r}(cos \alpha+i sin\alpha)---------(i)"

where "\alpha =(\theta +360^{\circ}k)/n,"

we can take k=0,1,2,3,..........,(n-1)

so using the above relation we can get the nth root of the complex number and which is in the form of z bar as "z= r(cos\u03b8+sin\u03b8 i)"

Now, take conjugate of general equation as

"\bar{z}=a-ib"

In polar coordinates form it will be "\bar {z} = r(cos \theta -i sin \theta )"

now again using De Moivre's theorem If we find roots in the conjugate case

"\sqrt [n]{\bar{z}}=\sqrt[n]{r}(cos \alpha-i \sin\alpha) -------(ii
)"

so from equation (i) and (ii) real parts of the nth part of the equation "\sqrt [n]{\bar{z}}=\sqrt [n]{{z}}" are the same.

Hence proved.